Calculate the H3O+1 concentration in 0.00500 M NaOH.
1
Expert's answer
2020-04-28T12:19:26-0400

Kw = 1*10-14 = [H3O+][OH-]

The amount of dissociated water is relatively small, so we count C(OH-) that was only formed by dissociation of NaOH.

[OH-] = 0.005 M

[H3O+] = Kw/[OH-] = 1*10-14/0.005 = 2*10-12